# Neural timescales from a computational perspective

### v1

@@ 4.1 Emergent timescales in task-performing neural networks
Computational neuroscience has had a long-standing interest in the repertoire of available timescales to elucidate how neural circuits can represent and perform cognitive tasks with complex temporal structures. Classically, models were “hand-crafted”, i.e., parameters were manually selected or analytically determined to satisfy a priori constraints, and were intended to capture how neuronal circuits implement dynamics necessary for computation, such as delayed persistent activity for working memory[96] or slow evidence accumulation for decision-making[97, 98]. To understand the underlying mechanisms, related theoretical investigations focused on how tailored structures in recurrent network connectivity and its eigenspectrum shape the dynamical landscape or phase portrait of the network[99, 100] (Box 2), with special interests in, e.g., fixed points, line attractors, and limit cycles (Fig. 4b), which could be leveraged for memory[101, 102], input integration[103, 100], and timing[104] respectively (see [100] for a recent review on attractor models of neural computation). Crucially, the timescales of the latent dynamical system near and on these low-dimensional manifolds are emergent properties of interest, since they govern, for example, how quickly and how stably a network can integrate new input while remembering or forgetting old ones, filter out fast noisy fluctuations to sense slower signals, and receive inputs of different frequencies.
@@
Concurrently, related research on reservoir computing, including liquid state machines [105] and echo state networks [106], similarly considered the activity timescales of RNNs from the perspective of latent dynamical systems, since they require long-term but not divergent (or exploding) dynamics at the edge-of-chaos [107, 108]. While reservoir networks typically have trained input and output weights, the recurrent weights are fixed due to the challenges of propagating error over many time steps, and much attention was dedicated to investigating how they can be thoughtfully initialized to maximize the network’s computational capacity (similar to [99]). Particularly relevant was how input-driven chaos in such networks can be tuned to represent a flexible repertoire of timescales (or temporal filters) for time-dependent computations [107].
[t] Box 2. Activity timescales of a linear network As an illustrative example, we review how to calculate the activity timescale ($\tau_{\textrm{fit}}$) of a linear recurrent network (i.e., linear dynamical system). Consider the system $\frac{dx(t)}{dt}=Ax(t)$ where $x(t)$ is the state vector, e.g., neural activity at time $t$, and $A$ is a square connectivity matrix. To acquire the timescales of the system dynamics, we solve for the eigenvalues of the matrix $A$, $\lambda_{i}$, through the characteristic equation, $A\textbf{u}_{i}=\lambda_{i}\textbf{u}_{i}$. For a stable decaying mode, i.e., non-exploding, the real part of the associated eigenvalue must be negative, and its decay timescale is defined as $\tau_{i}=-\frac{1}{\textrm{Re}(\lambda_{i})}.$ The dominant activity timescale of the system (i.e., $\tau_{\textrm{fit}}$) is then typically determined by the mode whose eigenvalue has the largest real part. Furthermore, for a linear network composed of units with fixed membrane time constant $\tau_{\textrm{param}}$, the system dynamics follow $\tau_{param}\frac{dx(t)}{dt}=Ax(t).$ Assuming $\tau_{\textrm{param}}$ to be the same for all neurons, the activity timescale is scaled accordingly, $\tau_{i}=-\frac{\tau_{\textrm{param}}}{\textrm{Re}(\lambda_{i})}.$ Extensions to networks with further biological constraints, e.g., random networks obeying Dale’s law, can also be found [99].
These earlier studies often considered task constraints from a theoretical and analytical perspective [109] (Fig. 4c, left), and they motivated recent investigations using task-optimized networks to tackle similar questions in more complex or analytically intractable networks. The advancement of modern machine learning hardware accelerated network optimization strategies for RNNs (e.g., backpropagation-through-time [110] via stochastic gradient descent) and enabled a complementary approach: instead of handcrafting networks with certain dynamical properties required for a task, the recurrent weights and other parameters of networks can be optimized directly to perform tasks or to mimic experimental recordings [111] (Fig. 4c, middle and right), and subsequently analyzed to identify the repertoire of neural dynamics important for the required computation (i.e., “opening the black box”[102]). Consistent with earlier works on hand-crafted networks, analyses often reveal fixed points corresponding to different inputs or memories [112], line attractors for tracking continuous environmental or behavioral variables [113], and (quasi-)limit cycles or state space trajectories traversed at different speeds that correspond to the timing of network inputs or outputs [49, 50], while multi-task networks exhibit context- and task phase-dependent composition of the above “computational primitives” [114, 115].
In particular, efficient optimization of computational models has opened up recent investigations into the connection between observed neural activity and the underlying latent dynamical system. By fitting models of neural population activity with latent states governed by, for example, switching linear dynamics, one could extract the time constants of dynamical objects of interest, such as approximate line attractors representing evidence or internal state accumulation [113]. Moreover, one could directly compare the temporal dynamics of individual neurons’ activity in experimental measurements with that of neurons in task-performing network models (Fig. 4d), both in rate-based and spiking recurrent neural networks, to dissect cellular, network, and learning-related mechanisms of computations underlying e.g., temporal scaling [49], working memory [81, 116], and other motor or cognitive tasks [117].
@@
These ideas can be further extended to study networks endowed with more sophisticated and biologically-realistic constraints, such as Dale’s law[118, 81] and dendritic compartmentalization [119], hypothesized neural computational mechanisms such as oscillatory phase coding[120, 121], as well as better analytical tractability (e.g., low-rank RNNs[122]) or improved expressiveness, computational power, and trainability [123, 124, 125]). Targeting the same sought-after dynamical repertoires, “hybrid” approaches can leverage more sophisticated weight initialization or regularization strategies to nudge the network toward line attractors or limit cycles after training [126, 125]. Furthermore, since population activity timescales reflect the state of neural dynamics relative to critical transitions, timescales have also been used to study how the state of dynamics in spiking neuromorphic networks affect the performance of tasks with varying complexity[127].
@@
Altogether, these studies provide insights into specific tasks or network architectures underlying computations of interest from the perspective of network dynamics. Leaning on the notion that the geometry of neural population representation is intimately linked to (the timescales of) its dynamics and that the latter may be even more important than the former [128], both classical and recent works lend support to the importance of the diversity of neural timescales for different computations. In other words, the temporal signatures of a dynamical system are a reflection of its phase portrait geometry, and constraining one to satisfy certain computational requirements often means constraining the other as well[124]. Nevertheless, it may be possible for networks to acquire similar low-dimensional geometries but vastly different timescales of dynamics. Therefore, we suggest explicit comparisons of in silico and in vivo timescales in task-performing networks as a straightforward yet synergistic research direction to provide explicit evidence for the necessity (or the lack thereof) of different neural timescales.
@@ 4.2 Optimization of timescale-related parameters
Another powerful way to use modern ANNs to study the effect of neural timescales on computation is to optimize timescale-related parameters in the network model, such as synaptic, membrane, and adaptation time constants, or more phenomenological variables, such as temporal receptive field size and context length. In a standard RNN, all neurons are given a single fixed time constant that is roughly analogous to the membrane time constant of biological neurons, which determines how quickly the effect of an input on that neuron decays. In reality, neurons across the brain have a range of membrane time constants due to varying properties such as dendritic morphology and the composition of conductances (see Section 3.1). As such, time constants can be individually optimized (similar to the connectivity weight matrices) [129, 130, 131, 132, 117], or even parameterized to be input-dependent, arriving at gated network architectures (e.g., long short-term memory units, or LSTMs). Thus, by comparing to networks that are not endowed with diverse timescale parameters, such studies provide evidence for the potential necessity of different timescales in behaviorally relevant computations.
@@
Several recent studies have shown that spiking networks trained to perform standard machine learning or cognitive tasks (e.g., delayed classification, working memory maintenance, etc.) can benefit from both the heterogeneity and explicit optimization of single-neuron or spike-frequency adaptation time constants [81, 130, 61, 131, 116]. Strikingly, the trained networks exhibit similar timescale distributions as membrane time constants[130] or activity timescales of individual neurons in analogous brain regions of animals performing the same task[81]. Furthermore, time constants do not necessarily need to match that of the task, but both longer and diverse time constants were found to be beneficial [61]. Similarly, RNNs with trainable time constants demonstrate improved task performance and memory capacity while the learned time constants mimic the hierarchy of timescales in the data-generating process[133, 134]. With more bio-realistic RNNs, such as those with connectivity based on the measured connectome of the fruit fly visual system, training both the connection weights and single-neuron time constants resulted in emergent stimulus tuning similar to that of in vivo recorded neurons in the fly[132].
@@
Remarkably, multi-layer feed-forward deep neural networks trained to simply mimic the input-output relationship of a single morphologically accurate pyramidal neuron recover complex and spatially varying temporal filters [135], further emphasizing the computational capacity provided by multiple timescales, even at the scale of single neurons. Recurrent architecture fitted to the same morphologically accurate pyramidal neuron allows for the explicit implementation of multiple timescales within a neuron’s model and enables a single neuron with capabilities to solve complex long temporal dependencies tasks [136]. On the other hand, diverse activity timescales can arise from network interactions instead of cellular properties (as discussed in the previous section). A recent study suggests that developing network-mediated diverse and long timescales may be a more robust mechanism for performing long memory tasks rather than trainable time constants [116]. Hence, future studies need to further distill the computational role of diverse cellular or network-mediated timescales.
@@
These principles extend to more complex deep learning-based network architectures, providing functional benefits on more difficult tasks. Instead of trainable and fixed time constants, they can be further made to be input-dependent via trained gating weights (similar to gated RNNs in DL, such as GRUs and LSTMs), which improves modeling of sequences with multiple timescales [137]. Furthermore, initializing gating parameters to reflect the distribution of timescales in the data improves performance on tasks involving memorizing and operating on long sequences [129]. Leveraging this observation, endowing LSTM units with a distribution of effective timescales constructed to match the power-law temporal dependency of natural language improves network performance in language modeling, especially long-range dependencies [138]. In parallel, these “chrono initialized” LSTMs improved decoding of brain response to narrative text, as well as estimation of voxel-wise fMRI BOLD timescales along the cortical hierarchy in human subjects[139], uniting potential principles of operation underlying both artificial and biological neural networks.
@@
Finally, recent research in deep learning has shown an intriguing convergence of principles based on thoughtful inductive biases and initialization, as well as self-supervised pre-training, for improving performance on tasks with long-range dependencies [140]. Although RNN-based architectures can model infinitely long temporal dependencies in theory (similar to IIR filters), simpler implementations based on finite-length temporal convolutions appeared to be more effective at acquiring long timescale memories [141], though both architectures learn effective temporal context windows that reflect natural structures in data (e.g., speech [142]). More recent architectures based on deep (linear) state space models (SSMs, e.g., S4 [143] and others [144, 145]) enjoy the benefits of both recurrence and convolutions, outperforming Transformer-based architectures on long timescale tasks, which are hindered by quadratic complexity in time. Later extensions capitalize on these insights to derive better parameterization or initialization strategies, resulting in competitive models with simpler implementations, such as Structure Global Convolution [146] and Linear Recurrence Units [147]. On the other hand, it has been shown that both Transformer-based and SSM-based architectures can be drastically improved through self-supervised pre-training on the input sequence data, prior to downstream task training (e.g., labeled classification) on the same long-range tasks [148]. These findings suggest that equipping deep learning models with inductive biases appropriate for long timescale tasks can be promising, but the exact mechanisms of these improvements—and whether they are related to intrinsic timescales in biological networks—remain unclear.
@@
In summary, there is a growing interest in convergent approaches to tuning network models to produce emergent task-relevant dynamics and timescales, as well as explicit optimization of time-constant parameters in ANNs. These approaches appear to be a promising and collaborative avenue for research in neuroscience and machine learning aiming to understand and improve time-dependent computations in biological and artificial systems.
@@ 5 Discussion
In this paper, we discussed how different computational methods help to turn empirical observation of diverse neural timescales into quantitative and testable hypotheses about brain dynamics: Different timescale estimation methods can disentangle various components of neural activity, such as modality- or task-specific variables, from intrinsic dynamics, while mechanistic models constrained by experimentally observed timescales, together with task-performing ANNs, can elucidate potential circuit mechanisms underlying brain dynamics and computations. Integrating these directions, we here provide an outlook for future research and discuss open questions.
@@ 5.2 Models and data, an inseparable loop
Computational models suggest that various cellular and network mechanisms could alter timescales in the brain. Therefore, timescale-constrained models can be used as a great tool to provide mechanistic insights into brain dynamics and computations. It is important to note again that activity timescales of neurons may differ significantly from time-constant parameters implemented in the models (e.g., membrane or synaptic time constants) (Box 1). Hence, it is crucial to match the activity timescales of the models to those of the data. For models with fewer mechanisms or free parameters, it is possible to derive the activity timescales analytically[70, 79]. However, in the brain, multiple mechanisms interact together to shape timescales, and studying their interactions analytically is not straightforward. In such cases, qualitative matching, e.g., generating a similar order of hierarchy of timescales[46, 72] or a grid search over multiple model parameters[62, 8] are used to fit the model timescales to data.
@@
5.3 Timescales seed further synergies between neuroscience and ML
@@
At the same time, modern ML models such as large language models or deep reinforcement learning models can facilitate studying the role of neural timescales in naturalistic tasks occurring over longer timescales than what can be studied in common laboratory-constrained experiments. Indeed, a recent study on freely behaving animals reported a different organization of neural timescales across brain areas than what has been found during resting state[9]. Therefore, timescales can serve as a key to open new avenues where ML tools can expedite uncovering mechanisms underlying brain computation.

### v2

@@ 4.1 Emergent timescales in task-performing neural networks
Computational neuroscience has had a long-standing interest in the repertoire of dynamics—and by extension, their timescales—to elucidate how neural circuits can represent and perform tasks with complex temporal structures. For instance, in classical hand-tuned networks, fixed points, line attractors, and limit cycles (Fig. 5c) are dynamical motifs that can be leveraged for memory [136, 137], input integration and decision-making [138, 139, 140], and timing [141], respectively (see [142] for a recent review on attractor models of neural computation). Beyond the geometry itself, timescales of the latent dynamical system near and on these low-dimensional manifolds are also key properties of interest. Crucially, those timescales govern, for example, how quickly and how stably a network can integrate new input while remembering or forgetting old ones, filter out fast, noisy fluctuations to sense slower signals, and receive inputs of different frequencies. More recently, and consistent with hand-tuned networks [143], task- and activity-optimized models also exhibit fixed points corresponding to persistent inputs or memories [144], line attractors for tracking slow environmental or behavioral variables [34, 35, 145, 146], and (quasi-)limit cycles traversed at different speeds corresponding to the timing of network inputs or outputs [9, 71].
@@
Additionally, related research on reservoir computing, including liquid state machines [147] and echo state networks [148], similarly considered the activity timescales of RNNs from the perspective of latent dynamical systems, since they require long-term but not divergent (or exploding) dynamics at the edge-of-chaos [149, 150]. While reservoir networks typically have trained input and output weights, the recurrent weights are fixed due to the challenges of propagating errors over many time steps. As such, much attention was dedicated to investigating how weights can be thoughtfully initialized to maximize the network’s computational capacity [151, 152], and in particular, how input-driven chaos in such networks can be tuned to represent a flexible repertoire of timescales (or temporal filters) for time-dependent computations [149].
Notably, an appropriate distribution of timescales is not only important for implementing these computations, they are useful for identifying key dynamical structures in the first place. For instance, an asymmetry in the decay time constants of a (locally) linear system (Box 3) is indicative of one slow direction and many fast directions—in other words, a line attractor [153, 34]. Identification of such behaviorally-relevant latent structures in network models further enables causal manipulation of the neuronal populations involved, closing the loop between theory and experiment in dissecting circuit mechanisms underlying complex behaviors such as the role of subcortical brain structures in mating and encoding of affect in self and others [35, 145, 146].
@@
Alternatively, one could directly compare the activity timescales of individual neurons in experimental recordings with that of neurons in task-performing SNNs and RNNs (Fig. 5d). This approach can similarly aid in understanding cellular, network, and learning-related mechanisms of computations underlying e.g., temporal scaling [9], working memory [114, 154], and other motor or cognitive tasks [155]. These ideas can be further extended to study networks endowed with more sophisticated and biologically-realistic constraints such as Dale’s law [156, 114] and dendritic compartmentalization [157], hypothesized computational mechanisms such as oscillatory phase coding [158, 159], as well as better analytical tractability (e.g., low-rank RNNs [160]), improved expressiveness, and trainability [161, 162, 163]. Targeting the same sought-after dynamical repertoires, “hybrid” approaches can leverage more sophisticated weight initialization or regularization strategies to nudge the network toward line attractors or limit cycles after training [164, 163]. Furthermore, since population activity timescales reflect the state of neural dynamics relative to critical transitions, timescales have also been used to study how the state of dynamics in spiking neuromorphic networks affect the performance of tasks with varying complexity [165].
@@
Altogether, these studies provide insights into specific tasks or network architectures underlying computations of interest from the perspective of cellular and network dynamics. Leaning on the notion that the geometry of representation in a neural population is intimately linked to (the timescales of) its dynamics, and that the latter may be even more important than the former [166], both classical and recent works lend support to the importance of diverse neural timescales for different computations. In other words, the temporal signatures of a dynamical system are a reflection of its phase portrait geometry, and constraining one to satisfy certain computational requirements often means constraining the other as well [162].
Nevertheless, it may be possible for networks to acquire similar low-dimensional geometries but vastly different timescales of dynamics. Therefore, we suggest explicit comparisons of in silico and in vivo timescales in task-performing networks as a straightforward yet synergistic research direction to provide explicit evidence for the necessity (or lack thereof) of different timescales of dynamics. It should be noted, however, that similar to nonlinear biophysical networks discussed in Section 3.2 and Box 3, analyses of timescales in the context of task-performing models similarly rest on the assumption of linearized systems in a particular state. While such analyses provide insights on “local” dynamics and computational properties, individual timescales do not inform of global properties of the network, which may require sampling a distribution of timescales across multiple states and task contexts.
@@ 4.2 Optimization of timescale-related parameters
Another powerful way to use modern ANNs to study the effect of neural timescales on computation is to directly optimize timescale-related parameters in the network model, such as synaptic, membrane, and adaptation time constants, or more phenomenological variables like temporal receptive field size and context length:
@@
In a standard RNN, all neurons are given a single fixed time constant that is roughly analogous to the membrane time constant of biological neurons, which determines how quickly the effect of an input on that neuron decays. In reality, neurons across the brain have a range of membrane time constants due to varying properties such as dendritic morphology and the composition of conductances (see Section 3.1). As such, time constants can be individually optimized as parameters (similar to connectivity weights) [167, 168, 169, 170, 155, 154], or even parameterized to be input-dependent in gated network architectures (e.g., long short-term memory units, or LSTMs). Thus, by comparing to networks not endowed with diverse time constants, such studies provide evidence for the potential necessity of different timescales in computations.
@@
For instance, several recent studies have shown that both recurrent and feedforward spiking networks trained to perform machine learning or cognitive tasks can benefit from both the heterogeneity and explicit optimization of single-neuron or spike-frequency adaptation time constants [114, 168, 84, 169, 171]. Strikingly, the trained networks exhibit similar timescale distributions as membrane time constants [168] or activity timescales of individual neurons in analogous brain regions of animals performing the same task [114]. Furthermore, time constants do not necessarily need to match that of the task, but both longer and diverse time constants were found to be beneficial [84]. Similarly, RNNs with trainable time constants demonstrate improved task performance and memory capacity while the learned time constants recover the hierarchy of timescales in the data-generating process [172, 173]. With more bio-realistic RNNs, such as those constrained by connectome measurements, training both the connection weights and single-neuron time constants resulted in emergent stimulus tuning similar to that of in vivo recorded neurons in the fly [170].
@@
Remarkably, multi-layer feed-forward deep neural networks trained to simply mimic the input-output relationship of a single morphologically accurate pyramidal neuron recover complex and spatially varying temporal filters [174], further emphasizing the computational capacity provided by multiple timescales, even at the scale of single neurons. Recurrent architecture fitted to the same morphologically accurate pyramidal neuron allows a single neuron to solve tasks with complex long temporal dependencies [175]. On the other hand, diverse activity timescales can arise from network interactions instead of cellular properties (as discussed in the previous section). A recent study suggests that developing network-mediated diverse and long timescales may be a more robust mechanism for performing long memory tasks rather than trainable time constants [154]. Hence, future studies need to further distill the computational role of diverse cellular or network-mediated timescales.
@@
These principles extend to more complex deep learning-based network architectures, providing functional benefits on more difficult tasks. Instead of trainable and fixed time constants, they can be further made to be input-dependent via trained gating weights (similar to gated RNNs, such as LSTMs and Gated Recurrent Units), which improves modeling of sequences with multiple timescales [176]. Furthermore, initializing gating parameters to reflect the distribution of timescales in the data improves performance on tasks involving memorizing and operating on long sequences [167]. Leveraging this observation, endowing LSTM units with a distribution of effective timescales constructed to match the power-law temporal dependency of natural language improves network performance in language modeling, especially long-range dependencies [177]. In parallel, these “chrono initialized” LSTMs improved decoding of brain response to narrative text, as well as estimation of voxel-wise fMRI BOLD timescales along the cortical hierarchy in human subjects [178], uniting potential principles of operation underlying both artificial and biological neural networks.
@@
Finally, recent research in deep learning has shown an intriguing convergence of principles based on thoughtful inductive biases and initialization, as well as self-supervised pretraining, for improving performance on tasks with very long temporal dependencies [179]. Although RNN-based architectures can model infinitely long temporal dependencies in theory (similar to IIR filters), simpler implementations based on finite-length temporal convolutions appeared to be more effective at acquiring long timescale memories [180], though both architectures learn effective temporal context windows that reflect natural structures in data (e.g., speech [181]).
More recent architectures based on deep linear state space models (SSMs, e.g., LMU [182], S4 [183], and others [184, 185, 186, 187]) enjoy the benefits of both recurrence and convolutions, outperforming Transformer-based architectures on long timescale tasks, which are hindered by quadratic complexity in time. Remarkably, the success of these linear SSM-based models was in part driven by replacing the standard dense recurrent connections with structured and/or diagonal complex matrices, which directly parameterize the eigenvalues of the linear system—in other words, the system timescales (Box 3). Later extensions capitalize on these insights to derive better parameterization or initialization strategies, resulting in competitive models with simpler implementations, such as Structure Global Convolution [188] and Linear Recurrence Units [186]. On the other hand, it has been shown that both Transformer-based and SSM-based architectures can be drastically improved through self-supervised pretraining on the input sequence data, prior to downstream task training (e.g., labeled classification) on the same long-range tasks [189]. These findings suggest that pretraining or directly equipping deep learning models with inductive biases appropriate for long timescale tasks can be promising, but the exact mechanisms of these improvements—and to what extent they are related to intrinsic timescales in biological networks—remain unclear.
@@
In summary, there is growing interest in convergent approaches to tuning network models to produce emergent task-relevant dynamics and timescales, as well as explicit optimization of time constant parameters in ANNs. Leveraging these approaches, findings from biological and deep learning network models suggest that both the magnitude and diversity of timescales may be important for tasks and data with long-range dependencies. Thus, these approaches appear to be a promising and collaborative avenue for research in neuroscience and AI aiming to understand and improve time-dependent computations in biological and artificial systems.
@@ 5 Discussion
We discussed how different computational methods help to turn empirical observation of diverse neural timescales into quantitative and testable hypotheses about brain dynamics: Different timescale estimation methods can disentangle various components of neural activity, such as modality- or task-specific variables, from intrinsic dynamics, while mechanistic models constrained by experimentally observed timescales, together with task-performing ANNs, can elucidate potential circuit mechanisms underlying brain dynamics and computations. Integrating these directions, we here provide an outlook for future research and discuss open questions.
@@ 5.2 Models and data, an inseparable loop
Computational models suggest that various cellular and network mechanisms could alter timescales in the brain. Therefore, timescale-constrained models can be used as a great tool to provide mechanistic insights into brain dynamics and computations. It is important to note again that activity timescales of neurons may differ significantly from time constant parameters implemented in the models (e.g., membrane or synaptic time constants) (Box 2). Hence, it is crucial to match the activity timescales of the models to those of the data. For models with fewer mechanisms or free parameters, it is possible to derive the activity timescales analytically [104, 112]. However, in the brain, multiple mechanisms interact together to shape timescales, and studying their interactions analytically is not straightforward. In such cases, qualitative matching, e.g., generating a similar order of hierarchy of timescales [68, 106] or a grid search over multiple model parameters [17, 24] are used to fit the model timescales to data.
@@
5.3 Timescales seed further synergies between neuroscience and AI
@@
At the same time, modern AI models, such as large language models or deep reinforcement learning models, can facilitate studying the role of neural timescales in naturalistic tasks occurring over longer timescales than what can be studied in common laboratory-constrained experiments. Indeed, a recent study on freely behaving animals reported a different organization of neural timescales across brain areas than what has been found during resting state [41]. Therefore, timescales can serve as a key to open new avenues where ML tools can expedite uncovering mechanisms underlying brain computation.